# Some Aspects of Three-Quark Potentials (Part II)

### v2

@@ IV.1.1 Configuration A
We are now ready to analyze the $-\ell$ behavior of $E^{\text{\,\tiny A}}$. The analysis is based on the observation that small $\ell$ corresponds to small $v$, as follows from the asymptotics (A.2) and (A.5). Taking the limit $v\rightarrow 0$ in Eqs.(4.1) and (4.3), we get
@@
$\eta=\frac{\cos^{-\frac{1}{2}}\alpha_{1}\bigl{[}1+I\bigl{(}\sin^{2}\alpha_{1},\tfrac{1}{2},\tfrac{3}{4}\bigr{)}\bigr{]}+\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}{\cos^{-\frac{1}{2}}\alpha_{2}\bigl{[}1+I\bigl{(}\sin^{2}\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr{)}\bigr{]}-\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}\,.$ (4.6)
At a given $\eta$, the solution to these equations determines the limiting values of the tangent angles. We then expand  Asuperscript $\ell^{\text{\,\tiny A}}$ and $E^{\text{\,\tiny A}}$ near $v=0$ and reduce the two equations to a single one, with the result
@@
In studying the $-\ell$ behavior, we take both $\lambda$’s to approach $1$. In that limit, the tangent angles of the side strings go to $\cos\alpha_{1}=\cos\alpha_{2}=v\text{e}^{1-v}$, so that $\alpha_{1}=\alpha_{2}$. Solving the force balance equations under this condition gives: $\sin\alpha_{1}=(3\mathsf{k}(1+4v)\text{e}^{-3v}-1)/2$ and $\alpha_{3}=\pi/2$. The compatibility of the two expressions for $\alpha_{1}$ implies that
@@
This is precisely Eq.(3.27) evaluated at $\theta=\pi$. Importantly, its solution ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is independent of $\eta$ and provides an upper bound on the parameter $v$, which thus varies from 00 to ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$.
The leading terms in the expansion of  Asuperscript $\ell^{\text{\,\tiny A}}$ and $E^{\text{\,\tiny A}}$ can be found using the asymptotics (A.4) and (A.10). So,
@@
Thus, the $-\ell$ expansion takes the form
@@ IV.1.2 Configuration B
We now briefly discuss configuration B. The only difference from configuration A is that the tangent angle $\alpha_{1}$ is non-negative. As a result, the corresponding expressions can be obtained from those of configuration A by replacing, for string 1, the functions ${\cal L}^{-}$ and ${\cal E}^{-}$ with ${\cal L}^{+}$ and ${\cal E}^{+}$. After doing so, we get
@@
As usual, when studying the $-\ell$ behavior, we consider the limit $v\rightarrow 0$. In this limit, the force balance equations reduce to (4.5), while the geometrical constraint becomes
$\eta=\frac{\cos^{-\frac{1}{2}}\alpha_{1}I\bigl{(}\cos^{2}\alpha_{1},\tfrac{3}{4},\tfrac{1}{2}\bigr{)}+\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}{\cos^{-\frac{1}{2}}\alpha_{2}\bigl{(}1+I\bigl{(}\sin^{2}\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr{)}\bigr{)}-\cos^{-\frac{1}{2}}\alpha_{3}\,I\bigl{(}\cos^{2}\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr{)}}\,.$ (4.16)
@@ IV.2 Some examples
[…]
@@ IV.3 A remark on diquarks
$L=\begin{cases}L^{\text{\,\tiny A}}(v)&\text{for}\quad\text{v}_{\ast}\leq v\leq{\text{v}}_{\text{\tiny 0}}\,,\\ L^{\text{\,\tiny B}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,,\end{cases}\qquad E_{\text{\tiny 3Q}}=\begin{cases}E^{\,\text{\tiny A}}(v)&\text{for}\quad\text{v}_{\ast}\leq v\leq{\text{v}}_{\text{\tiny 0}}\,,\\ E^{\,\text{\tiny B}}(v)&\text{for}\quad{\text{v}}_{\text{\tiny 0}}\leq v\leq{\text{v}}_{\text{\tiny 1}}\,\,,\end{cases}$ (4.24)
@@
To proceed further, we consider $\ell$ to be small enough assuming that the diquark $[Q_{1}Q_{3}]$ can be regarded as nearly point-like. Our first goal is to examine the three-quark potential as a function of $L$. The corresponding plots are shown in the left panel of Figure 8. As seen, $E_{\text{\tiny 3Q}}$ behaves linearly at large $L$ and becomes nearly Coulomb-like
Figure 8: $E_{\text{\tiny 3Q}}$(left) and $\delta$ (right) as a function of $L$
@@
[…]
@@ V.1 The leading correction to the $Y$-law
Figure 9: A string configuration relevant in the IR limit, with the negative tangent angles. The arrows on the $xy$-plane indicate the vectors $\vec{\text{e}}$.
@@
A string becomes infinitely long as its $\lambda$ parameter approaches $1$, as explained in Appendices A and B. In this limit, the tangent angle at the baryon vertex is given by $\cos\alpha=v\text{e}^{1-v}$, and it is the same for all the strings shown in the Figure. Because of this, the vectors $\vec{\text{e}}$ all have the same length. This implies that the angles between them must be $\frac{2}{3}\pi$ to satisfy the first equation in (5.1), and therefore, the point $Y$ coincides with the Fermat point $F$ of the triangle. [^17] Note that this argument doesn’t extend to triangles with angles greater or equal to $\frac{2}{3}\pi$, because in such cases, $F$ coincides with one of the triangle’s vertices, and the corresponding string can’t be infinitely long. Thus, we recover the $Y$-law. Moreover, from the second equation in (5.1), we find that $\sin\alpha=\mathsf{k}(1+4v)\text{e}^{-3v}$. Combining this with the above expression for $\cos\alpha$ leads directly to Eq.(3.12) whose solution is ${\text{v}}_{\text{\tiny 1}}$. The above derivation of the $Y$-law is more physical than that of a3Q2008 , where it was derived by solving a variational problem.
@@
For our parameters set, the right hand side is positive, with its value given in (3.35). This implies that $3c>C_{\text{\tiny 3Q}}$, or, in other words, the constant term in the IR limit is smaller than that in the UV limit.
@@ V.2 What happens when a triangle has an angle greater than $\frac{2}{3}\pi$?
where $\text{v}={\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ for the plus sign and $-\text{v}={\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ for the minus sign. In the last step, we used the relation $\gamma_{1}=\pi-\frac{\theta_{3}}{2}$. An interesting observation here is that the right hand side depends only on the obtuse angle $\theta_{3}$, and not on $\ell$ or $L$. The result for collinear geometry in Sec.IV is a special case of this general behavior.
So far we have considered the IR limit without referring to the parameter values. More specifically, as seen in Figure 23, the solution $-{\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ exists only in a tiny interval near $\theta_{3}=\frac{2}{3}\pi$. This implies that outside this interval, the IR limit is governed by a single string configuration with $\alpha_{3}>0$. The potential thus behaves as
@@ VI Conclusions
a composite object dressed by gluons, as in the Figure on the right, and due to the epsilon tensor transforms in the two-index antisymmetric representation. As we saw in Sec.V, when the largest angle exceeds $\frac{2}{3}\pi$, one of the quarks does approach the baryon vertex in the IR limit, while the other recede. Using $[Q]$, the leading term in Eq.(5.15) can be immediately recovered as coming from two strings stretched between the quark sources. Notably, a similar construction is possible with an antiquark. In the context of tetraquarks, a $[Q][\bar{Q}]$ pair called the $Q\bar{Q}$ seeding was suggested in isgur and recently found in the current model aQQqq .
@@
(3) The issue of the subtracted potential komas can be addressed within the current model. The key point is that the constant terms differ in the UV and IR limits. [^19] This discrepancy was first found in the case of the equilateral triangle geometry a3Q2016 . If, therefore, $E_{\text{\tiny 3Q}}-\sigma L_{\text{min}}-3c$ does not vanish in the IR limit, $E_{\text{\tiny 3Q}}-\sigma L_{\text{min}}-C_{\text{\tiny 3Q}}$ does. As seen from Figure 17 of komas , the minimum discrepancy (in absolute value) arises in the collinear geometry, while the maximum in several geometries where the largest angle does not exceed $\frac{2}{3}\pi$ [^20] Due to limited data, it is not clear what happens with the obtuse-wide geometry considered in komas . Our model predicts that its discrepancy value lies between the two extrema. This agrees with the result shown in Figure 11 as the discrepancy originates from the difference $C_{\text{\tiny 3Q}}-3c$. There is another important implication. We saw in Sec.IV that in the case of the collinear geometry, the potential $E_{\text{\tiny 3Q}}$ deviates from the $\Delta$-law. The deviation arises because the difference $C_{\text{\tiny 3Q}}-\frac{3}{2}C_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is non-zero. Note that such a difference identically vanishes in the UV limit, where $03c-\frac{3}{2}\cdot 2c=$ 0.
@@ Acknowledgements.
We would like to thank P. de Forcrand, Z. Komargodski, M. Shifman, P.Weisz, and S. Zhong for useful discussions and correspondence concerning this topic. This work was conducted as part of Program FFWR-2024-0012 at the Landau Institute.
@@ Appendix A Notation and useful formulas
${\cal L}^{+}$ is a non-negative function that vanishes when $\alpha=\frac{\pi}{2}$ or $x=0$. It exhibits a singularity at $(0,1)$. If $\alpha$ approaches $\alpha_{0}$ as $x$ approaches 00, then the behavior of ${\cal L}^{+}$ for small $x$ is
${\cal L}^{+}(\alpha,x)=\sqrt{x}\bigl{(}{\cal L}^{+}_{0}(\alpha_{0})+O(x)\bigr{)}\,,\qquad\text{with}\qquad{\cal L}^{+}_{0}(\alpha_{0})=\frac{1}{4}\cos^{-\frac{1}{2}}\alpha_{0}\,B\bigl{(}\cos^{2}\alpha_{0};\tfrac{3}{4},\tfrac{1}{2}\bigr{)}\,.$ (A.2)
@@
[…]
@@
${\cal E}^{+}(\alpha,x)=\frac{1}{\sqrt{x}}\int^{1}_{0}\,\frac{du}{u^{2}}\,\biggl{(}\text{e}^{xu^{2}}\Bigl{[}1-\cos^{2}{}\alpha\,u^{4}\text{e}^{2x(1-u^{2})}\Bigr{]}^{-\frac{1}{2}}-1-u^{2}\biggr{)}\,,\qquad 0\leq\alpha\leq\frac{\pi}{2}\,,\qquad 0\leq x\leq 1\,.$ (A.7)
The function ${\cal E}^{+}$ is singular at $x=0$ and at the point $(0,1)$. If $\alpha$ approaches $\alpha_{0}$ as $x\rightarrow 0$, then the small-$x$ behavior of ${\cal E}^{+}$ is
${\cal E}^{+}(\alpha,x)=\frac{1}{\sqrt{x}}\bigl{(}{\cal E}^{+}_{0}(\alpha_{0})+O(x)\bigr{)}\,,\qquad\text{with}\qquad{\cal E}_{0}^{+}(\alpha_{0})=\frac{1}{4}\cos^{\frac{1}{2}}\alpha_{0}\,B\bigl{(}\cos^{2}\alpha_{0};-\tfrac{1}{4},\tfrac{1}{2}\bigr{)}\,;$ (A.8)
@@
The function ${\cal E}^{-}$ is singular at $(0,0)$ and at $y=1$. More precisely, as $y$ approaches $1$ with $x$ held fixed, it behaves like
@@
Numerically, $I_{0}=0.751$. This function is increasing and vanishes at $x=0.278$. It is related to the functions ${\cal L}^{-}$ and ${\cal E}^{-}$ via

### v3

@@ IV.1.1 Configuration A
We are now ready to analyze the small-$\ell$ behavior of $E^{\text{\,\tiny A}}$. The analysis is based on the observation that small $\ell$ corresponds to small $v$, as follows from the asymptotics (A.2) and (A.5). Taking the limit $v\rightarrow 0$ in Eqs.(4.1) and (4.3), we get
@@
$\eta=\frac{\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{1}\bigl[1+I\bigl(\sin^{2}\hskip-2.56073pt\alpha_{1},\tfrac{1}{2},\tfrac{3}{4}\bigr)\bigr]+\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}{\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{2}\bigl[1+I\bigl(\sin^{2}\hskip-2.56073pt\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr)\bigr]-\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}\,.$ (4.6)
At a given $\eta$, the solution to these equations determines the limiting values of the tangent angles. We then expand  $\ell^{\text{\,\tiny A}}$ and $E^{\text{\,\tiny A}}$ near $v=0$ and reduce the two equations to a single one, with the result
@@
In studying the large-$\ell$ behavior, we take both $\lambda$’s to approach $1$. In that limit, the tangent angles of the side strings go to $\cos\alpha_{1}=\cos\alpha_{2}=v\text{e}^{1-v}$, so that $\alpha_{1}=\alpha_{2}$. Solving the force balance equations under this condition gives: $\sin\alpha_{1}=(3\mathsf{k}(1+4v)\text{e}^{-3v}-1)/2$ and $\alpha_{3}=\pi/2$. The compatibility of the two expressions for $\alpha_{1}$ implies that
@@
This is precisely Eq.(3.27) evaluated at $\theta=\pi$. Importantly, its solution ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is independent of $\eta$ and provides an upper bound on the parameter $v$, which thus varies from 0 to ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$.
The leading terms in the expansion of  $\ell^{\text{\,\tiny A}}$ and $E^{\text{\,\tiny A}}$ can be found using the asymptotics (A.4) and (A.10). So,
@@
Thus, the large-$\ell$ expansion takes the form
@@ IV.1.2 Configuration B
[…]
@@
As usual, when studying the small-$\ell$ behavior, we consider the limit $v\rightarrow 0$. In this limit, the force balance equations reduce to (4.5), while the geometrical constraint becomes
$\eta=\frac{\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{1}I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{1},\tfrac{3}{4},\tfrac{1}{2}\bigr)+\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}{\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{2}\bigl(1+I\bigl(\sin^{2}\hskip-2.56073pt\alpha_{2},\frac{1}{2},\tfrac{3}{4}\bigr)\bigr)-\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{3}\,I\bigl(\cos^{2}\hskip-2.56073pt\alpha_{3},\tfrac{3}{4},\frac{1}{2}\bigr)}\,.$ (4.16)
@@ IV.2 Some examples
where $C_{\text{\tiny 3Q}}=C^{\,\text{\tiny A}}$. Notably, the $\Delta$-law yields the same leading term. Since it also provides a good approximation at small $\ell$, one may expect that the difference between the genuine three-quark potential and the $\Delta$-law is small. To explore this, we compare the two in the right panel of the Figure, which plots the potential derived from the $\Delta$-law. The visible difference arises from the constant terms in the large-$\ell$ expansions and amounts to approximately $57\,\text{MeV}$, as indicated in Table II. This difference could potentially be determined with reasonable accuracy through numerical simulations.
@@ IV.3 A remark on diquarks
[…]
@@
To proceed further, we consider $\ell$ to be small enough assuming that the diquark $Q_{1}Q_{3}]$ can be regarded as nearly point-like. Our first goal is to examine the three-quark potential as a function of $L$. The corresponding plots are shown in the left panel of Figure 8. As seen, $E_{\text{\tiny 3Q}}$ behaves linearly at large $L$ and becomes nearly Coulomb-like
Figure 8: $E_{\text{\tiny 3Q}}$(left) and $\delta$ (right) as a function of $L$ .
@@
as motivated by heavy quark-diquark symmetry. [^15] Alternatively, the constant term in (4.26) can be fixed by imposing $\delta=0$ as $L\rightarrow\infty$. Importantly, $\delta$ doesn’t depend on the normalization constant $c$. Our results for $\delta$ are shown in the right panel of the Figure as a function of $L/\ell$. As expected, $\delta$ decreases with increasing $L$. Quantitatively, the value of $\delta$ drops below $3\,\text{MeV}$ for $L\gtrsim 4\ell$, suggesting that the diquark approximation is reasonable for phenomenological applications in the nonperturbative region. [^16] However, it was reported that the accuracy of the diquark approximation diminishes with increasing temperature bakry .
@@ V.1 The leading correction to the $Y$-law
Figure 9: A string configuration relevant in the IR limit, with the negative tangent angles. The arrows on the xyxy-plane indicate the vectors $\vec{\text{e}}$.
@@
A string becomes infinitely long as its $\lambda$ parameter approaches $1$, as explained in Appendices A and B. In this limit, the tangent angle at the baryon vertex is given by $\cos\alpha=v\text{e}^{1-v}$, and it is the same for all the strings shown in the Figure. Because of this, the vectors $\vec{\text{e}}$ all have the same length. This implies that the angles between them must be $\frac{2}{3}\pi$ to satisfy the first equation in (5.1), and therefore, the point $Y$ coincides with the Fermat point $F$ of the triangle. [^17] Note that this argument doesn’t extend to triangles with angles greater or equal to $\frac{2}{3}\pi$, because in such cases, $F$ coincides with one of the triangle’s vertices, and the corresponding string cannot be infinitely long. Thus, we recover the $Y$-law. Moreover, from the second equation in (5.1), we find that $\sin\alpha=\mathsf{k}(1+4v)\text{e}^{-3v}$. Combining this with the above expression for $\cos\alpha$ leads directly to Eq.(3.12) whose solution is ${\text{v}}_{\text{\tiny 1}}$. The above derivation of the $Y$-law is more physical than that of a3Q2008 , where it was derived by solving a variational problem.
@@
For our parameter set, the right hand side is positive, with its value given in (3.35). This implies that $3c>C_{\text{\tiny 3Q}}$, or, in other words, the constant term in the IR limit is smaller than that in the UV limit.
@@ V.2 What happens when a triangle has an angle greater than $\frac{2}{3}\pi$?
where $\text{v}={\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ for the plus sign and $\text{v}={\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ for the minus sign. In the last step, we used the relation $\gamma_{1}=\pi-\frac{\theta_{3}}{2}$. An interesting observation here is that the right hand side depends only on the obtuse angle $\theta_{3}$, and not on $\ell$ or $L$. The result for collinear geometry in Sec.IV is a special case of this general behavior.
So far we have considered the IR limit without referring to the parameter values. More specifically, as seen in Figure 23, the solution ${\text{v}}_{\text{\tiny 1}}^{\text{\small-}}$ exists only in a tiny interval near $\theta_{3}=\frac{2}{3}\pi$. This implies that outside this interval, the IR limit is governed by a single string configuration with $\alpha_{3}>0$. The potential thus behaves as
@@ VI Conclusions
a composite object dressed by gluons, as in the Figure on the right, and due to the epsilon tensor transforms in the two-index antisymmetric representation. As we saw in Sec.V, when the largest angle exceeds $\frac{2}{3}\pi$, one of the quarks does approach the baryon vertex in the IR limit, while the other recede. Using $[Q]$, the leading term in Eq.(5.15) can be immediately recovered as coming from two strings stretched between the quark sources. Notably, a similar construction is possible with an antiquark. In the context of tetraquarks, a $Q][\bar{Q}]$ pair called the $Q\bar{Q}$ seeding was suggested in isgur and recently found in the current model aQQqq .
@@
(3) The issue of the subtracted potential komas can be addressed within the current model. The key point is that the constant terms differ in the UV and IR limits. [^19] This discrepancy was first found in the case of the equilateral triangle geometry a3Q2016 . If, therefore, $E_{\text{\tiny 3Q}}-\sigma L_{\text{min}}-3c$ does not vanish in the IR limit, $E_{\text{\tiny 3Q}}-\sigma L_{\text{min}}-C_{\text{\tiny 3Q}}$ does. As seen from Figure 17 of komas , the minimum discrepancy (in absolute value) arises in the collinear geometry, while the maximum in several geometries where the largest angle does not exceed $\frac{2}{3}\pi$. [^20] Due to limited data, it is not clear what happens with the obtuse-wide geometry considered in komas . Our model predicts that its discrepancy value lies between the two extrema. This agrees with the result shown in Figure 11 as the discrepancy originates from the difference $C_{\text{\tiny 3Q}}-3c$. There is another important implication. We saw in Sec.IV that in the case of the collinear geometry, the potential $E_{\text{\tiny 3Q}}$ deviates from the $\Delta$-law. The deviation arises because the difference $C_{\text{\tiny 3Q}}-\frac{3}{2}C_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is non-zero. Note that such a difference identically vanishes in the UV limit, where $3c-\frac{3}{2}\cdot 2c=0$.
@@ Acknowledgements.
We would like to thank P. de Forcrand, Z. Komargodski, M. Shifman, P.Weisz, and S. Zhong for useful discussions and correspondence concerning this topic. This work was conducted as part of Program FFWR-2024-0011 at the Landau Institute.
@@ Appendix A Notation and useful formulas
${\cal L}^{+}(\alpha,x)=\cos\alpha\sqrt{x}\int^{1}_{0}du\,u^{2}\,\text{e}^{x(1-u^{2})}\Bigl[1-\cos^{2}{}\hskip-2.84526pt\alpha\,u^{4}\text{e}^{2x(1-u^{2})}\Bigr]^{-\frac{1}{2}}\,,\qquad 0\leq\alpha\leq\frac{\pi}{2}\,,\qquad 0\leq x\leq 1\,.$ (A.1)
@@
${\cal L}^{+}(\alpha,x)=\sqrt{x}\bigl({\cal L}^{+}_{0}(\alpha_{0})+O(x)\bigr)\,,\qquad\text{with}\qquad{\cal L}^{+}_{0}(\alpha_{0})=\frac{1}{4}\cos^{-\frac{1}{2}}\hskip-2.56073pt\alpha_{0}\,B\bigl(\cos^{2}\hskip-2.56073pt\alpha_{0};\tfrac{3}{4},\tfrac{1}{2}\bigr)\,.$ (A.2)
@@
${\cal L}^{-}(y,x)=\sqrt{y}\biggl(\,\int^{1}_{0}du\,u^{2}\,\text{e}^{y(1-u^{2})}\Bigl[1-u^{4}\,\text{e}^{2y(1-u^{2})}\Bigr]^{-\frac{1}{2}}+\int^{1}_{\sqrt{\frac{x}{y}}}du\,u^{2}\,\text{e}^{y(1-u^{2})}\Bigl[1-u^{4}\,\text{e}^{2y(1-u^{2})}\Bigr]^{-\frac{1}{2}}\,\biggr)\,,\quad 0\leq x\leq y\leq 1\,.$ (A.3)
@@
${\cal E}^{+}(\alpha,x)=\frac{1}{\sqrt{x}}\int^{1}_{0}\,\frac{du}{u^{2}}\,\biggl(\text{e}^{xu^{2}}\Bigl[1-\cos^{2}{}\hskip-2.84526pt\alpha\,u^{4}\text{e}^{2x(1-u^{2})}\Bigr]^{-\frac{1}{2}}-1-u^{2}\biggr)\,,\qquad 0\leq\alpha\leq\frac{\pi}{2}\,,\qquad 0\leq x\leq 1\,.$ (A.7)
[…]
${\cal E}^{+}(\alpha,x)=\frac{1}{\sqrt{x}}\bigl({\cal E}^{+}_{0}(\alpha_{0})+O(x)\bigr)\,,\qquad\text{with}\qquad{\cal E}_{0}^{+}(\alpha_{0})=\frac{1}{4}\cos^{\frac{1}{2}}\hskip-2.56073pt\alpha_{0}\,B\bigl(\cos^{2}\hskip-2.56073pt\alpha_{0};-\tfrac{1}{4},\tfrac{1}{2}\bigr)\,;$ (A.8)
@@
[…]
@@
[…]